OpenStudy (anonymous):

3x^2-11x-20/25-x^2 simplify

OpenStudy (mertsj):

\[\frac{(3x+4)(x-5)}{(5-x)(5-x)}\]

OpenStudy (mertsj):

Note that the quotient of opposites is -1. So cancel out x-5 and 5-x and get

OpenStudy (mertsj):

\[\frac{-(3x+4)}{(5-x)}\]

OpenStudy (anonymous):

so 3x+4/5-x is the answer ?

OpenStudy (mertsj):

Did you see the negative sign in front of the quantity 3x+4?

OpenStudy (mertsj):

This is the answer in 2 forms:

OpenStudy (mertsj):

\[\frac{-(3x+4)}{(5-x)} = \frac{(3x+4)}{(x-5)}\]

OpenStudy (mertsj):

Revised answer:\[\frac{-(3x+4)}{5+x}\]